OpenStudy (anonymous):

How do I simplify the expression (-5+i)/2i

OpenStudy (anonymous):

is it\[(-5+i)/2i\]

OpenStudy (anonymous):

Yeah

OpenStudy (anonymous):

divide is the word you are looking for or "write in standard form as \(a+bi\)

OpenStudy (anonymous):

\[\frac{-5+i}{2i}\times \frac{i}{i}\] is a start

OpenStudy (anonymous):

you get \[\frac{-5i-1}{-2}\]as a first step

OpenStudy (anonymous):

Okay, I've gotten that far... What about after that?

OpenStudy (anonymous):

then in standard form it would be \[\frac{1}{2}+\frac{5}{2}i\]

OpenStudy (anonymous):

How do you get there from the last step?

OpenStudy (anonymous):

break it in to two parts

OpenStudy (anonymous):

\[\frac{-5i-1}{-2}=\frac{-1}{-2}+\frac{-5}{-2}i=\frac{1}{2}+\frac{5}{2}i\]

OpenStudy (anonymous):

Oh, I see. Sorry, I just didn't see what was done between those two. So is that the final step?

OpenStudy (anonymous):

yeh

OpenStudy (anonymous):

Okay. Thanks a lot everyone!

OpenStudy (anonymous):

yw:)